The separation between two 1 kg masses is (a) decreased by 2/3, and (b) increased by a factor of 3. How is the mutual gravitational force affected in each case?
Answer by Alan3354(69443)

The separation between two 1 kg masses is (a) decreased by 2/3, and (b) increased by a factor of 3. How is the mutual gravitational force affected in each case?
--------------
Attraction is a function of the square of 1/d, --> k/d^2
-----
a) decreased by 2/3
--> 1/(2/3)^2 = 1/(4/9) = 9/4
increases by 9/4, or 2.25 as much attraction.
==============
and (b) increased by a factor of 3.
(1/3)^2 = 1/9
1/9 the attraction
======================
Notice the masses are not relevant.